OpenStudy (anonymous):

Solve Phase Plane equation dx/dt =y-1 dy/dt=e^(x+y)

OpenStudy (anonymous):

dy/dx = e^x e^y -------- y-1 (y-1)/e^y dy= e^x dx

OpenStudy (anonymous):

-e^(-y) y= e^x +C C= e^x +e^(-y) y